An iron railroad rail is 790 ft long when the temperature is 40°C. What is its length (in ft) when the temperature is −17°C? (Round your answer to at least one decimal place.)
Answer #1

Given,

The length of an iron rail road, L0 = 790 ft

Temperature T1 = 40 oC and T2 = -17 oC

The coefficient of thermal expansion of iron, a = 12 * 10-6 /C

We know,

L = Lo ( 1 + a \DeltaT)

= 790 (1 + 12 * 10-6 (40 + 17))

= 790 (1 + 12 * 10-6 (57))

= 790.54 ft